A kicked football leaves the ground at an angle θ = 45⁰ with the horizontal has an initial speed of 25 m/s. Determine the distance of X. Acceleration due to gravity is 10 m/s².
1
Expert's answer
2020-11-10T09:59:56-0500

Let's first find the time that the football takes to reach the maximum height from the kinematic equation:

"v_y=v_0sin\theta-gt_{rise},""0=v_0sin\theta-gt_{rise},""t_{rise}=\dfrac{v_0sin\theta}{g}."


Then, we can find the total flight time of the football:


"t=2t_{rise}=\dfrac{2v_0sin\theta}{g}."

Finally, we can find the horizontal range of the football from the kinematic equation:


"x=v_0tcos\theta=v_0cos\theta\dfrac{2v_0sin\theta}{g}=\dfrac{v_0^2sin2\theta}{g},""x=\dfrac{(25\ \dfrac{m}{s})^2\cdot sin2\cdot 45^{\circ}}{10\ \dfrac{m}{s^2}}=62.5\ m."

Answer:

"x=62.5\ m."